# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 13 Prospects of improving bounds on the interference terms
$\mathbf{E}_{G\in\mathcal{G}_{n\times n}}(|\mathsf{Per}(G)|^{2t})=O\bigg{(}\frac{(n!)^{2t}(t!)^{2n}}{(nt)!}\bigg{)}.$ (49)
@@
where $X_{i}$ are as defined in eqn. (43), and are independent, identically distributed random variables, $i\in\{1,\dots,N\}$. It is immediate to observe that $\mathbf{E}(Y_{i})=0$. We show the following in appendix 22.
@@
$N\frac{\mathbf{E}(|Y|^{r})}{\sigma_{N}^{r}}\geq O\Big{(}\frac{\beta(r)^{n}}{N^{\frac{r}{2}-1}}\Big{)},$ (51)
where $\beta(r)>1$ is a positive real number dependent on $r$.
When $N=\mathsf{poly}(n)$, Thm. 17 shows that Lyapunov condition is not satisfied. Although this condition is sufficient, but not necessary, for the CLT to hold we provide numerical evidence that $N=n^{2}$, $N=n^{3}$, and $N=n^{4}$ sized sums of i.i.d. Gaussian matrices do not converge to a normal distribution for $n\in\{2,3,4,5\}$. We plot our results for $N=n^{3}$ and $n\in\{2,3,4,5\}$ in Figures 6 (a)-(d).
It is interesting to note that in Thm. 17, when $N=O(\mathsf{exp}(n))$, the lower bound on the Lyapunov condition can converge to 0 as $n\to\infty$. Marginal probabilities corresponding to a large number of lost photons are sums of, possibly exponential, numbers of probabilities having the form of eqn. (43). Our result provides evidence that these marginals are asymptotically normally distributed, and therefore efficient to sample from. Although low order as well as high loss marginals of boson sampling are known to be easy to compute and sample from [58, 67, 17], our result might provide a new perspective on simulating lossy boson sampling marginals by linking these to normally distributed random variables.

### v3

@@ 16.1 Prospects of improving bounds on the interference terms
$\mathbf{E}_{G\in\mathcal{G}_{n\times n}}(|\mathsf{Per}(G)|^{2t})\in\Theta\bigg(\frac{(n!)^{2t}(t!)^{2n}}{(nt)!}\bigg).$ (96)
@@
where $X_{i}$ are as defined in eqn. (90), and are independent, identically distributed random variables, $i\in\{1,\dots,N\}$. It is immediate to observe that $\mathbf{E}(Y_{i})=0$. We show the following.
@@
$N\frac{\mathbf{E}(|Y|^{r})}{\sigma_{N}^{r}}\geq A\frac{\beta(r)^{n}}{N^{\frac{r}{2}-1}},$ (98)
where $\beta(r)>1$ is a positive real number dependent on $r$, and $A>0$ a constant.
When $N=\mathsf{poly}(n)$, Thm. 31 shows that Lyapunov condition is not satisfied. Although this condition is sufficient, but not necessary, for the CLT to hold we provide numerical evidence that $N=n^{2}$, $N=n^{3}$, and $N=n^{4}$ sized sums of i.i.d. Gaussian matrices do not converge to a normal distribution for $n\in\{2,3,4,5\}$. We plot our results for $N=n^{3}$ and $n\in\{2,3,4,5\}$ in Figures 8 (a)-(d).
It is interesting to note that in Thm. 31, when $N\in O(\mathsf{exp}(n))$, the lower bound on the Lyapunov condition can converge to 0 as $n\to\infty$. Marginal probabilities corresponding to a large number of lost photons are sums of, possibly exponential, numbers of probabilities having the form of eqn. (90). Our result provides evidence that these marginals are asymptotically normally distributed, and therefore efficient to sample from. Although low order as well as high loss marginals of boson sampling are known to be easy to compute and sample from [49, 69, 3], our result might provide a new perspective on simulating lossy boson sampling marginals by linking these to normally distributed random variables.